OpenStudy (anonymous):

What polynomial has roots of –4, 2, and 5 ?

OpenStudy (anonymous):

x3 – x2 – 22x + 40 x3 + x2 – 22x – 40 x3 + 3x2 – 18x – 40 x3 – 3x2 – 18x + 40

OpenStudy (anonymous):

try 1?

jimthompson5910 (jim_thompson5910):

x = -4 or x = 2 or x = 5 x+4=0 or x-2=0 or x-5=0 (x+4)(x-2)(x-5) = 0 See where to go from here?